OpenStudy (anonymous):

the quotient of a number and four is twelve less than the number what is the number

OpenStudy (kirbykirby):

\[\frac{a}{4}=a-12\]?

OpenStudy (kirbykirby):

a=4a-48 3a=48 a=16